OpenStudy (littleman789):

What is the equation in point-slope form of the line passing through (3, 6) and (−2, 1)?

OpenStudy (wckeller):

First use the two points in the slope formula \[m=\frac{ x_2-x_1 }{ y_2-y_1 }\] This will get you the slope for point-slope form. \[m=\frac{-2-3 }{ 1-6 }\] \[m=\frac{ -5 }{ -5 }\] \[m=1\] Now use one of the points and plug it into the point-slope form equation. \[y-y_1=m(x-x_1)\] \[y-6=1(x-3)\] Then solve for Y \[y=x+3\]